Find the solution of the system of equations below using substition.
y=3x+5
2x-y=-7

Answer by mananth(16946)

y=3x+5
2x-y=-7
plug the value of y in the second equation
2x-(3x+5)=-7
2x-3x-5=-7
.
-x==-2
x=2
.
y=11